# Breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov–Ivanov equation

### v1

@@ 4 Double-pole breather-to-soliton transitions
$K_{1}$ $=32t\cdot(2\beta_{1}+1)^{2}(2\beta_{1}-1)^{2}(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{3},$
$\quad\cdot\left(4i\beta_{1}^{2}-i-\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right),$
$K_{2}$ $=32\sqrt{2}\sqrt{-\beta_{1}^{8}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)}$
$\quad\cdot\sqrt{(\alpha_{1}^{2}+\beta_{1}^{2})^{3}t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}},$
$K_{3}$ $=\frac{\beta_{1}^{4}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)(\alpha_{1}^{2}+\beta_{1}^{2})\sqrt{2}}{2\sqrt{-\beta_{1}^{8}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)}}$
$\quad\cdot\frac{1}{\sqrt{(\alpha_{1}^{2}+\beta_{1}^{2})^{3}t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}}},$
$K_{4}$ $=\frac{t(2\beta_{1}+1)(2\beta_{1}-1)(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{2}}{24}$
$\quad\cdot\left(4i\beta_{1}^{2}-i-\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right).$
$\bullet$  When $e^{D_{2}}\sim t$, $D_{2}\to-\infty$,
@@
$L_{1}$ $=32t(2\beta_{1}+1)^{2}(2\beta_{1}-1)^{2}(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{3}$
$\quad\cdot\left(4i\beta_{1}^{2}-i+\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right),$
$L_{2}$ $=32\sqrt{2}\sqrt{t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{3}\beta_{1}^{8}}$
$\quad\cdot\sqrt{\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i-2\beta_{1}^{2}+2\alpha_{1}^{2}+1\right)},$
$L_{3}$ $=-\frac{t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{2}\beta_{1}^{4}\sqrt{2}}{\sqrt{t^{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{3}\beta_{1}^{8}}},$
$\quad\cdot\frac{1}{\sqrt{\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}i-2\beta_{1}^{2}+2\alpha_{1}^{2}+1\right)}},$
$L_{4}$ $=\frac{t(2\beta_{1}+1)(2\beta_{1}-1)(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{2}}{24}$
$\quad\cdot\left(4i\beta_{1}^{2}-i+\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right).$
We can find that the characteristic lines of the asymptotic solitons above are $\frac{D_{2}\Delta}{8}+\ln K_{3}=0$ and $\frac{D_{2}\Delta}{8}+\ln L_{3}=0$, both contain the logarithmic terms which can reflect the double-pole properties. When $a=\frac{3}{10},c=1,\delta=-\frac{3}{20},\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i}$ are taken, we can find that the characteristic trajectories of the corresponding asymptotic solitons in Fig.14 fit perfectly with the contour plot of Fig.13. Furthermore, we compare the asymptotic solitons (7), (8) and the exact solution (5) when the parameters are given. We can find that the asymptotic soliton perfectly matches the exact solution in the far-field region as is shown in Fig.15. This demonstrates the validity of our asymptotic analysis.
Figure 14: The contour plot for the double-pole anti-dark soliton.
(a) $t=30$
(b) $t=-30$
(c) $t=60$
(d) $t=-60$
Figure 15: (a)The comparison of the asymptotic soliton and the exact solution (black line) for the double-pole anti-dark soliton. (a)(c) Red: the asymptotic soliton $[u]_{\mathrm{I}}^{+}$; Green: the asymptotic soliton $[u]_{\mathrm{II}}^{+}$; (b)(d) Red: the asymptotic soliton $[u]_{\mathrm{I}}^{-}$; Green: the asymptotic soliton $[u]_{\mathrm{II}}^{-}$.
@@ 5 Conclusions and discussions
In conclusion, we explore the breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov–Ivanov equation.
$\bullet$  We obtain different types of the nonlinear converted waves, such as the W-shaped soliton, M-shaped soliton, multi-peak soliton, anti-dark soliton and periodic wave solution.
$\bullet$  The interactions between various noninear converted waves are exhibited.
$\bullet$  The double-pole breather-to-soliton transitions are studied and the asymptotic analysis method is applied to analyze the asymptotic behavior of the double-pole anti-dark soliton.
Our work will provide a reference for future projects about the breather-to-soliton transitions.
6 Data availability
No data was used for the research described in the article.

### v2

@@ 4 Double-pole breather-to-soliton transitions
$K_{1}$ $=32(2\beta_{1}+1)^{2}(2\beta_{1}-1)^{2}(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{3}$
$\quad\times\left(4\mathrm{i}\beta_{1}^{2}-\mathrm{i}-\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right)t,$
$K_{2}$ $=32\sqrt{2}\sqrt{-\beta_{1}^{8}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\mathrm{i}+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)}$
$\quad\times\sqrt{(\alpha_{1}^{2}+\beta_{1}^{2})^{3}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}}|t|,$
$K_{3}$ $=\frac{\sqrt{2}\beta_{1}^{4}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\mathrm{i}+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)(\alpha_{1}^{2}+\beta_{1}^{2})}{2\sqrt{-\beta_{1}^{8}\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\mathrm{i}+2\beta_{1}^{2}-2\alpha_{1}^{2}-1\right)}}$
$\quad\times\frac{1}{\sqrt{(\alpha_{1}^{2}+\beta_{1}^{2})^{3}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}}|t|},$
$K_{4}$ $=\frac{(2\beta_{1}+1)(2\beta_{1}-1)(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{2}}{24}$
$\quad\times\left(4\mathrm{i}\beta_{1}^{2}-\mathrm{i}-\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right)t.$
$\bullet$  When $e^{-D_{2}}\sim t$,
@@
$L_{1}$ $=32(2\beta_{1}+1)^{2}(2\beta_{1}-1)^{2}(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{3}$
$\quad\times\left(4\mathrm{i}\beta_{1}^{2}-\mathrm{i}+\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right)t,$
$L_{2}$ $=32\sqrt{2}\sqrt{(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{3}\beta_{1}^{8}}$
$\quad\times\sqrt{\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\mathrm{i}-2\beta_{1}^{2}+2\alpha_{1}^{2}+1\right)}|t|,$
$L_{3}$ $=-\frac{\sqrt{2}(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{2}\beta_{1}^{4}}{\sqrt{(2\beta_{1}+1)^{3}(2\beta_{1}-1)^{3}(4\alpha_{1}^{2}+1)^{3}(\alpha_{1}^{2}+\beta_{1}^{2})^{3}\beta_{1}^{8}}}$
$\quad\times\frac{|t|}{\sqrt{\left(\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\mathrm{i}-2\beta_{1}^{2}+2\alpha_{1}^{2}+1\right)}},$
$L_{4}$ $=\frac{(2\beta_{1}+1)(2\beta_{1}-1)(4\alpha_{1}^{2}+1)^{2}(\alpha_{1}^{2}+\beta_{1}^{2})\beta_{1}^{2}}{24}$
$\quad\times\left(4\mathrm{i}\beta_{1}^{2}-\mathrm{i}+\sqrt{(4\beta_{1}^{2}-1)(4\alpha_{1}^{2}+1)}\right)t.$
We can find the characteristic lines of the asymptotic solitons above are
$\frac{D_{2}\Delta}{8}+\ln K_{3}=0,~\frac{D_{2}\Delta}{8}+\ln L_{3}=0,$
both contain the logarithmic terms which can reflect the double-pole properties. When $a=\frac{3}{10},d=1,\delta=-\frac{3}{20},\lambda_{1}=\frac{7}{10}+\frac{\sqrt{39}}{10}\mathrm{i}$ are taken, it can be observed that the wave crest line graph of the corresponding asymptotic solitons in Fig.14 is associated with the contour plot in Fig.13. Furthermore, we compare the asymptotic solitons (12), (13) with the exact solution (11) under specific parameter settings. As is evident in Fig.15, the asymptotic solitons match the exact solution perfectly in the far-field region, which further validates the reliability of our asymptotic analysis.
Figure 14: The wave crest line graph (red line) for the asymtotic solitons of the double-pole anti-dark soliton in Fig.13 with $\alpha_{1}=\frac{7}{10},\beta_{1}=\frac{\sqrt{39}}{10}$. Auxiliary lines: Green: $t=0$; Black: 4x7.5752t=04x-7.5752t=0.
(a) t=5t=-5
(b) t=30t=-30
(c) t=60t=-60
(d) $t=5$
(e) $t=30$
(f) $t=60$
Figure 15: The comparison of the asymptotic solitons and the exact solution (green line) for the double-pole anti-dark soliton in Fig.13. (a)-(c) Red: the asymptotic soliton $u]_{\mathrm{I}}^{-}$; Black: the asymptotic soliton $u]_{\mathrm{II}}^{-}$; (d)-(f) Red: the asymptotic soliton $u]_{\mathrm{I}}^{+}$; Black: the asymptotic soliton $u]_{\mathrm{II}}^{+}$.
@@ 5 Conclusions and discussions
In conclusion, we explored the breather-to-soliton transitions and nonlinear wave interactions for the higher-order generalized Gerdjikov–Ivanov equation. The following key conclusions have been drawn:
$\bullet$  A rich family of the breather-to-soliton nonlinear converted waves, such as the W-shaped soliton, M-shaped soliton, multi-peak soliton, anti-dark soliton and periodic wave solution are obtained. Notably, these structures exhibit distinct dynamical features which have not been reported in lower-order GI models.
$\bullet$  The elastic interactions between various noninear converted waves were exhibited. These elastic interactions are characterized by the fact that the amplitude, velocity, and shape of the localized waves remain unchanged both before and after collision.
$\bullet$  The double-pole breather-to-soliton transitions were studied. Asymptotic analysis further demonstrated that the double-pole anti-dark solitons evolve into stable multi-peak structures under long-time propagation, a behavior directly attributed to the interplay between self-steepening and Raman terms in the HMGI equation.
Our work enriches the localized wave structures and the related dynamical properties for the HMGI equation. It will provide a reference for future projects about the breather-to-soliton transitions for the higher-order nonlinear evolution equations.
Acknowledgments
The authors thanks Prof. Q. P. Liu for helpful and enlightening suggestions. The work was supported by the Fundamental Research Funds for the Central Universities (No.3072025CFJ2405).
Data availability
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Author Declarations
The authors have no conflicts to disclose.